Grade 12IIT JEE Entrance Exam

The resultant of two vector u and v is perpendicular to the vector u and it`s magnitude is half of the magnitude of vector v angle between u and v

Rishu singh
9 Years agoGrade 12

1 Answer

SASANK reddy
9 Years ago
Ubar vector and Rbar vectors are perpendicular to each other.Rbar is resultant of Vbar and Ubar.magnitude of Rbar is V/2.angle between Vbar and Rbar is 30degrees.because cos(teta)=1/2.cos(teta)=v/2/V.so angle between Ubar and Vbar is 120 degrees